A 19.8 kg block is dragged over a rough, hor izontal surface by a constant force of 174 N acting at an angle of 25.8° above the horizon tal . The block is displaced 63.9 m, and the coefficient of kinetic friction is 0.123 .


1
Expert's answer
2021-11-16T19:19:15-0500

A 19.8 kg block is dragged over a rough, hor izontal surface by a constant force of 174 N acting at an angle of 25.8° above the horizon tal . The block is displaced 63.9 m, and the coefficient of kinetic friction is 0.123 .

a. Find the work done by the 174 N force. The acceleration of gravity is 9.8 m/s^2

b. Find the magnitude of the work done by the force of friction.

a.

"Work = Force \times Distance \times cos θ \\

Work = 174 \times 63.9 \times cos(25.8°) \\

Work = 10010.28 \;J"

b.

"N+F_y = P \\

N = P-F_y \\

N = m \times g -174 \times sin θ \\

N = 19.8 \times 9.8 -174 \times sin(25.8°) \\

N = 11897.88309 \;Newtons \\

F_a = N \times \mu \\

F_a = 118.309 \times 0.123 = 14.552 \; Newtons \\

Work = 14.552 \times 63.9 = 929.873 \;J"